OpenStudy (anonymous):

(h+9)^2=3 solve the equation

OpenStudy (anonymous):

you will get some numbers that do not "look very good" first take the square root of both sides: (h+9) = +-√(3) then subtract 9 from both sides h = √(3) - 9 h = -√(3) - 9

OpenStudy (anonymous):

h= 3 and -3 ??

OpenStudy (anonymous):

h = \[\sqrt{3}-9\] and \[-\sqrt{3}-9\]

OpenStudy (anonymous):

ohh ok thank you

OpenStudy (anonymous):

as I mentioned the answers don't "look pretty" but that's only because √3 is a non-repeating decimal.